# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: CSURccs: Computing methodologies Neural networksccs: Computer systems organization Distributed architecturesccs: Computer systems organization Data flow architecturesccs: Hardware Integrated circuits
@@ 1. Introduction
Graphs are a natural and powerful data structure for representing entities and their relationships. Many real-world data can be represented as graphs with nodes denoting a collection of entities and edges denoting their pairwise relationships, such as individuals in social networks, financial transactions between firms and banks, atoms and bonds in molecules, and vehicles in transportation systems. Graph neural networks (GNNs) (Kipf and Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017a) have recently become the most widely used graph machine learning (ML) model for learning knowledge and making predictions on graph data. GNNs have achieved state-of-the-art performance in many graph ML applications. They are used, for example, in recommendations on social graphs (Zhang et al., 2019a; Mao et al., 2021; Wu et al., 2020b), fraud account detection on financial graphs (Dou et al., 2020), drug discoveries from molecule graphs (Jiang et al., 2021), traffic forecasting on transportation graphs (Jiang and Luo, 2022), and so on.
@@
The superior performance of GNNs on graphs is mainly due to their ability to combine the entity information, represented as the node features, and the relationships, represented as the graph structure. GNNs take node features and the graph structure as input and output node representations, which can be used for various graph ML tasks like node classification, edge prediction, and graph classification. For each node in a graph, a GNN learns its representation by message passing, which aggregates messages (features) from the node’s neighbors and updates the representation by transforming the aggregated messages with a neural network. Each aggregate-and-update operation is called a GNN layer, which allows the node representation to combine information from the node’s direct neighbors. Stacking multiple GNN layers will apply such operations recursively so the node representation can gain information from multi-hop neighbors. Thus, the representation can capture the contextual information of a local neighborhood and be used to answer complex ML questions on graphs. For a target node with an $L$-layer GNN applied on it, all of its neighbors within $L$ hops form the receptive field of the target node. A recursively constructed $L$-layer tree graph with the target node as the root and the neighbors of each node as its children is called a computation graph of the target node.
@@
As the field of graph ML quickly develops, graph data is becoming gigantic with numerous nodes and edges in a single graph. For example, the Twitter user graph has 288M monthly active users (nodes) and an estimated 208 follow relations (edges) per user as of 3/2015, and the Facebook user graph has 1.39B active users and more than 400B total edges as of 12/2014 (Ching et al., 2015). These numbers have grown much larger and become hard to estimate. The commonly used academic benchmarks have also increased from thousands of nodes, e.g., Cora (Sen et al., 2008a), to as many as 240M nodes in the recent Open Graph Benchmark (OGB) (Hu et al., 2021). Large graphs raise scalability challenges for both GNN training and inference. Although deep neural networks (DNNs) for other data types also face scalability challenges when the number of DNN parameters becomes large, the challenge for GNNs is unique and there is greater emphasis on the large graph data (Zhang et al., 2020a; Jia et al., 2020b) rather than the large number of model parameters. When training DNNs on images or text, even if the amount of data is large, data can be processed by randomly sampling mini-batches because all data instances are assumed to be independent and identically distributed (iid). The DNN inference can also be done independently for each instance. For GNNs on graphs, however, the nodes are dependent on each other, which brings both benefits and challenges. GNNs leverage node dependency to learn informative representations and make predictions, but node dependency also means multi-hop neighbors are involved in the GNN computation for every single node. Therefore, it is nontrivial to partition graph nodes into mini-batches for GNN training considering the node dependency. Moreover, node dependency means that the sizes of computation graphs grow exponentially in the number of GNN layers (Alon and Yahav, 2021). On large graphs where deeper GNNs are often needed for better performance (Chen et al., 2020a; Li et al., 2021), not only does the mini-batch training need to process a huge computation graph but so does the inference on a single node. Take the Twitter user graph as an example, with 208 neighbors per node, a computation graph of a three-layer GNN involves all neighbors of a node within three steps, which can contain millions of nodes (i.e., roughly $208^{3}$ minus overlapped nodes). This node dependency problem makes applying GNNs on large graphs very challenging. A different situation for large datasets is a graph database with many graphs, for example, chemical molecules. Since graphs in these databases often only have dozens to hundreds nodes, the node dependency problem is not a big concern, and they are not the focus of this survey.
@@
Scalability challenges also show up for GNN systems, including both Commercial-off-the-Shelf (COTS) systems and customized hardware. Large computation graphs need to be processed because of the node dependency problem, but memory consumption becomes infeasible when the graph structure is represented as a huge adjacency matrix. In reality, the graph structure often needs to be processed in the sparse matrix format. However, sparse matrices have different computation patterns from standard DNNs. Therefore, the driving force of the success of deep learning on large data, i.e., the parallelism of computation devices like GPUs, cannot be fully exploited for efficient computation. This demands efficient systems with accelerated computation kernels and customized hardware accelerators like FPGA.

### v2

@@ Abstract.
copyright: ccjournal: CSURjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3805796ccs: Computing methodologies Neural networksccs: Computer systems organization Distributed architecturesccs: Computer systems organization Data flow architecturesccs: Hardware Integrated circuits
@@ 1. Introduction
Graphs are a natural and powerful data structure for representing entities and their relationships. Many real-world data can be represented as graphs with nodes denoting entities and edges denoting their pairwise relationships, such as individuals in social networks, financial transactions, atoms and bonds in molecules, and vehicles in transportation systems. Graph neural networks (GNNs) (Kipf and Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017b) have become the most widely used graph machine learning (ML) model on graph data. GNNs have achieved state-of-the-art performance in many applications, including recommendations on social graphs (Zhang et al., 2019b; Mao et al., 2021; Wu et al., 2020a), fraud detection on financial graphs (Dou et al., 2020), drug discovery from molecule graphs (Jiang et al., 2021), and traffic forecasting on transportation graphs (Jiang and Luo, 2022).
@@
The superior performance of GNNs stems from combining entity information (node features) and relationships (graph structure). GNNs take node features and graph structure as input and output node representations for various tasks like node classification, edge prediction, and graph classification. For each node, a GNN learns its representation through message passing, which aggregates messages (features) from the node’s neighbors and updates the representation via a neural network transformation. Each aggregate-and-update operation is called a GNN layer, allowing the representation to combine information from direct neighbors. Stacking multiple GNN layers applies such operations recursively, enabling the representation to capture information from multi-hop neighbors and the neighborhood context. For a target node with an $L$-layer GNN, all neighbors within $L$ hops form its receptive field. A recursively constructed $L$-layer tree graph with the target node as root and each node’s neighbors as children is called a computation graph.
@@
As graph ML develops rapidly, graph data has become gigantic. For example, the Twitter user graph had 288M monthly active users (nodes) and an estimated 208 follow relations (edges) per user as of 3/2015, while the Facebook user graph had 1.39B active users and over 400B total edges as of 12/2014 (Ching et al., 2015). Academic benchmarks have also grown from thousands of nodes, e.g., Cora (Sen et al., 2008), to as many as 240M nodes in the recent Open Graph Benchmark (OGB) (Hu et al., 2021). Large graphs raise scalability challenges for both GNN training and inference. Unlike DNNs for images or text, where data can be processed via random mini-batch sampling due to the iid assumption, GNN nodes are interdependent. While GNNs leverage this dependency to learn informative representations, it also means multi-hop neighbors are involved in computation for every single node. This makes partitioning graph nodes into mini-batches nontrivial and causes computation graphs to grow exponentially with the number of GNN layers (Alon and Yahav, 2021). On large graphs where deeper GNNs are often needed (Chen et al., 2020a; Li et al., 2021), both mini-batch training and single-node inference must process huge computation graphs. For instance, on the Twitter user graph with 208 neighbors per node, a three-layer GNN computation graph can contain millions of nodes (roughly $208^{3}$ minus overlaps). This node dependency problem makes applying GNNs on large graphs very challenging. Note that graph databases with many small graphs (e.g., chemical molecules with dozens to hundreds of nodes) do not face this issue and are not the focus of this survey.
@@
Scalability challenges also affect GNN systems, including both Commercial-off-the-Shelf (COTS) systems and customized hardware. Large computation graphs require substantial memory, making adjacency matrix representations infeasible. In practice, graph structures must be processed in sparse matrix format. However, sparse matrices have different computation patterns from standard DNNs, preventing full exploitation of parallelism on devices like GPUs. This demands efficient systems with accelerated computation kernels and customized hardware accelerators like FPGAs.